Performance Products Corporation makes two products, titanium Rims and Posts. Data regarding the two products follow:

   

Direct
Labor-Hours
per unit
Annual
Production
Rims 0.40 20,000 units
Posts 0.20 80,000 units

Additional information about the company follows:

  1. Rims require $17 in direct materials per unit, and Posts require $10.
  2. The direct labor wage rate is $16 per hour.
  3. Rims are more complex to manufacture than Posts and they require special equipment.
  4. The ABC system has the following activity cost pools:
Estimated Activity
Activity Cost Pool Activity Measure Estimated
Overhead
Cost
Rims Posts Total
Machine setups Number of setups $ 21,600 100 80 180
Special processing Machine-hours $ 180,000 4,000 0 4,000
General factory Direct labor-hours $ 288,000 8,000 16,000 24,000

2. Determine the unit product cost of each product according to the ABC system. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

unit product cost of rims: ?

Unit product cost of posts: ?
Answer #1

Answer

  • Requirement 2

Unit product cost of Rims

$                       37.00

Unit product cost of Posts

$                       15.72

  • Workings

Overhead cost

Activity

Activity rates

[A]

[B]

[C = A/B]

Machine Setup

$21,600

                                           180

$120.00

per setup

Special Processing

$180,000

                                       4,000

$45.00

per MH

General factory

$288,000

                                     24,000

$12.00

per DLHs

Manufacturing Overhead cost:

Rims

Posts

Machine Setup

$12,000

100 x $ 120

$9,600

80 x $ 120

Special Processing

$180,000

4000 x $ 45

$0

0 x $ 45

General factory

$96,000

8000 x $ 12

$192,000

16000 x $ 12

A

Total Overhead cost allocated

$288,000

$201,600

B

Total Units

                        20,000

                                 80,000

C = A/B

Overhead cost per unit

$                       14.40

$                                   2.52

D

Direct Material

$                       17.00

$                                10.00

E

Direct Labor Cost

$                         5.60

[0.4 hrs x $ 16]

$                                   3.20

[0.2 hrs x $ 16]

F = C+D+E

Answer: Unit Cost

$                       37.00

$                                15.72